Five years ago, Madge was five times as old as my youngest sister Veronica, Today I知 only three times as old as she is. How old is Madge?
Answer by ankor@dixie-net.com(22740)

Five years ago, Madge was five times as old as my youngest sister Veronica,
Today I知 only three times as old as she is. How old is Madge?
:
I think Madge and I are the same person (otherwise it doesn't make sense)
:
Let m = M's present age
Let v = V's present age
:
"Five years ago, Madge was five times as old as my youngest sister Veronica,
m - 5 = 5(v-5)
m - 5 = 5v - 25
m = 5v - 25 + 5
m = 5v - 20
:
Today I知 only three times as old as she is.
m = 3v
:
Replace m with 3v in the 1st equation
3v = 5v - 20
20 = 5v - 3v
20 = 2v
v = 10 is V's age
:
How old is Madge?
3(10) = 30 yrs is M's age